OpenStudy (anonymous):

f(x)=4cosx+7tanx find f'(x)

OpenStudy (anonymous):

the derivatives are "known" d/dx cos(x) = -sin(x) d/dx tan(x) = sec^2(x) so f'(x) = -4sin(x) + 7sec^2(x)

OpenStudy (anonymous):

wat happens to the (7tanx) and (4cosx)? @Euler271

OpenStudy (anonymous):

i get (-4sinx)(7tanx)+(4cosx)(7sec^2x)

OpenStudy (anonymous):

that is product rule. it is when you have F(x) = f(x)g(x), then F'(x) = f'(x)g(x) + g'(x)f(x) like you just did. this is simply F(x) = f(x) + g(x). derive seperately